# Many Minds from One Model: Bayesian Transformers for Population Intelligence

### v1

@@ Abstract
Despite their scale and success, modern transformers are almost universally trained as single-minded systems: optimization produces one deterministic set of parameters, representing a single functional hypothesis about the data. Motivated by the idea that intelligence emerge from many minds, we propose Population Bayesian Transformers (B-Trans), which transform a standard Large Language Model into a Bayesian Transformer model to supports sampling diverse yet coherent model instances from a single set of pre-trained weights.
B-Trans introduces a Bayesian-motivated posterior proxy by treating the bias-like offsets in normalization layers as stochastic variables with a Gaussian variational approximation, inducing a distribution over model behavior without the cost of training full Bayesian neural networks. Sampling from this proxy yields a set of model instances with diverse behaviors while maintaining general competence. To preserve coherence within each generation, we freeze the sampled noise at the sequence level, enforcing temporal consistency across tokens. B-Trans allows for population-level decision-making, where aggregating predictions across sampled individuals significantly enhances exploration. Experiments across zero-shot generation, Reinforcement Learning with Verifiable Rewards (RLVR), and RL without explicit labels demonstrate that B-Trans effectively leverage the wisdom of crowds, yielding superior semantic diversity while achieving better task performance compared to deterministic baselines.
@@
Bayesian Transformers for Population Intelligence
@@ 1 Introduction
The predominant paradigm in Large Language Model (LLM) yang2025qwen3; liu2025deepseek deployment treats model weights as static, deterministic point estimates. While this approach effectively leverages the knowledge acquired during pre-training wang2024mmlu, it inherently limits the model’s ability to represent uncertainty. Recent studies further suggest that well-aligned models suffer from severe structural homogenization (jiang2025artificial). Under the convergence pressures of post-training alignment, such as Supervised Fine-Tuning (SFT) and Reinforcement Learning from Human Feedback (RLHF), models tend to collapse onto a narrow set of reasoning patterns. This rigidity becomes a fundamental bottleneck in complex reasoning tasks: the model struggles to explore alternative reasoning paths, locking itself into local optima yue2025does.
@@
Current attempts to reintroduce diversity often operate at decoding time through randomness in the action space, for example via temperature-based sampling or similar entropy-promoting strategies. While increasing temperature can prevent verbatim repetition, it often acts as a surface-level perturbation. As noted by yue2025does, in reasoning tasks such sampling can lead to incoherent degeneration, yielding statistical variation without meaningful semantic progress, rather than systematically discovering novel solution paths. The core issue is that manipulating the output tokens does not alter the underlying deterministic reasoning engine. To recover diversity of thought, we argue that we must step back from a collapsed point estimate and allow the model to explore a local proxy of the posterior neighborhood around $\theta_{\text{MAP}}$, instead of relying solely on output-token randomization.
@@
Can we achieve meaningful population diversity without abandoning the efficiency of a single model? In this work, we demonstrate the feasibility by generating many minds that can emerge from a single model. Our key insight is that normalization layers ba2016layer; zhang2019root modulate representation geometry, and that small shifts in their bias-like offsets can induce meaningfully different functional behaviors while preserving architectural stability. We therefore replace point-estimated normalization biases with a Gaussian distribution, transforming a single transformer into a compact generator of a diverse population of model instances. Each draw from this distribution corresponds to a distinct “mind” that shares the same weights and structure, yet processes information through a different normalization regime. This targeted population modeling yields diversity where it is influential, without incurring the cost of maintaining multiple full models.
@@
To maintain logical coherence, we enforce temporal consistency by freezing the sampled offsets for the duration of each sequence. Unlike per-token noise injection (e.g., standard Dropout), which can lead to disjointed logic, our approach ensures that each generation is produced by a single, logically consistent “persona” sampled from the population. We present B-Trans as a general population formulation (Section 3) and instantiate it through normalization-layer perturbations (Section 4); this instantiation is deliberately minimal and can be extended in future work.
@@
We validate B-Trans through three distinct experimental lenses, each probing a different facet of population intelligence. First, in zero-shot generation, we demonstrate that B-Trans effectively breaks structural homogenization, producing functionally heterogeneous outputs. Second, in Reinforcement Learning with Verifiable Rewards (RLVR) guo2025deepseek, we show that the parameter-space diversity of B-Trans enables deep exploration, traversing sparse reward landscapes more effectively than action-space baselines. Finally, in a label-free Test-Time RL (TTRL) setting, we reveal that the implicit population utilizes the wisdom of crowds. Even without ground-truth supervision, aggregating the diverse reasoning paths of the B-Trans population outperforms deterministic baselines, showing that uncertainty can act not as noise, but as a generative engine for exploration. Collectively, these results confirm that B-Trans fundamentally alters the exploratory dynamics of the model, enabling deep and population-driven reasoning at negligible computational cost. We summarize our contribution as:
@@ 2 Related Works
Bayesian methods provide a principled framework for modeling weight uncertainty (blundell2015weight). Prior work has adapted Bayesian principles to neural networks for specific pragmatic objectives. For instance, kristiadi2020being demonstrated that being “a bit Bayesian” (e.g., applying Laplace approximation only to the last layer) can improve calibration and reduce overconfidence (“knowing what you don’t know”). Similarly, jing2025kalman introduced the Kalman Bayesian Transformer to support stable sequential learning and mitigate catastrophic forgetting by treating pre-trained weights as priors. While these works primarily use Bayesian tools for calibration, robustness, uncertainty quantification, or model averaging for prediction quality, B-Trans targets a generation-centric goal: we use lightweight parameter-space stochasticity to promote exploration and semantic diversity during generation. Compared to Deep Ensembles (lakshminarayanan2017simple) which scale linearly in cost, or full BNNs which are computationally intractable, B-Trans induces this population diversity solely through stochastic perturbations of normalization biases, offering a negligible alternative specifically optimized for the autoregressive nature of Transformers.
@@
Standard decoding strategies, such as Top-$k$ and Nucleus Sampling, primarily manipulate the tail of the token distribution to mitigate degeneration, but they do not address convergence in the underlying model. Recent work points out that modern instruction-tuned models exhibit severe structural homogenization, where the probability mass over-concentrates on a narrow set of reasoning patterns (jiang2025artificial). Consequently, purely stochastic decoding methods often yield variations of a single persona rather than distinct cognitive perspectives. Methods such as Self-Consistency wang2022self aggregate multiple sampled outputs, but remain constrained by the same model-level collapse. B-Trans instead diversifies behavior by sampling model instances via parameter-space perturbations, turning a single set of weights into a population that can produce more functionally diverse generations.
@@
Efficient exploration is critical for policy improvement, especially in reasoning tasks with sparse rewards. However, the efficacy of standard action-space entropy regularization (e.g., temperature sampling) is severely compromised by the aforementioned structural homogenization.
Recent RLVR studies observe that simply increasing output entropy often forces the model into incoherent degeneration (statistical noise) rather than systematically discovering alternative solution paths yue2025does. To address this, we revisit the concept of Parameter Space Noise plappert2018parameter, originally proposed for continuous control, and adapt it to modern LLMs. By perturbing a lightweight posterior proxy through B-Trans, we encourage deep exploration over reasoning trajectories beyond surface-level token randomization.
3 Bayesian Transformer
This section formalizes a population view of a single pre-trained model via Bayesian hypothesis sampling: we draw a sequence-level parameter sample to obtain one coherent model instance, and aggregate multiple individual for population-level thinking. [^1] Throughout, “Bayesian” is used as a principled motivation for structured uncertainty and sampling, rather than a claim of exact posterior inference for billion-parameter LLMs. Section 4 then presents a minimal instantiation that perturbs normalization offsets as a computationally negligible proxy for posterior sampling.
@@ 3.1 Local Geometry and the Laplace Approximation
Let $\mathcal{D}$ be the training dataset and $\theta\in\mathbb{R}^{d}$ be the parameters of a Transformer. We view the pre-trained parameters $\theta_{\text{MAP}}$ as the Maximum A Posteriori (MAP) estimate of the posterior $p(\theta|\mathcal{D})$. While $\theta_{\text{MAP}}$ captures a high-probability configuration, it collapses the rich information contained in the geometry of the loss landscape into a single point. To recover uncertainty, we employ the Laplace Approximation, modeling the posterior neighborhood around the mode as a Gaussian:
@@
Computing the exact covariance $\Sigma$ (the inverse Hessian) is not feasible for billion-parameter models. Furthermore, the true posterior landscape of LLMs is highly non-isotropic (li2018visualizing), and estimating its anisotropy at inference time without training data or gradients is not well-posed. However, explicitly modeling this anisotropy without access to the full training data or gradients during inference is ill-posed. For tractability, we adopt an isotropic covariance structure $\Sigma=\sigma^{2}I$. From an information-theoretic perspective, this choice represents the Maximum Entropy distribution given a constraint on the expected Euclidean distance from the $\theta_{\text{MAP}}$ estimate. Rather than risking a mis-specified diagonal approximation that might bias the model into degenerate subspaces, the isotropic assumption effectively defines a spherical trust region for exploration. Importantly, our goal is not exact Bayesian inference, but as efficient sampling from a locally valid proxy posterior, but a practical balancing theoretical motivation with engineering feasibility.
@@
A key challenge for Bayesian sampling in autoregressive generation is handling the time dimension. Standard stochastic regularizers, such as Dropout (srivastava2014dropout), are typically applied independently at each forward pass. For a sequence $y_{1:T}$, this implies integrating over the parameters at every token step:
@@
We argue that Eq. (2) is misaligned with coherent multi-step reasoning: it re-samples $\theta_{t}$ at each token, which can disrupt cross-token logical consistency by changing the underlying computation mid-generation. Instead, we adopt hypothesis sampling. In the Bayesian view, a single sample $\theta\sim p(\theta|\mathcal{D})$ corresponds to one coherent hypothesis. Once sampled, this hypothesis should remain fixed for the entire sequence. The predictive distribution is:
@@
We approximate Eq. (3) using Monte Carlo integration with $K$ samples. For each sample $k$, we draw a noise vector $\epsilon^{(k)}$ once and freeze it, obtaining a specific model member $\theta^{(k)}=\theta_{\text{MAP}}+\epsilon^{(k)}$. This ensures that the entire sequence $y_{1:T}$ is generated by a single consistent instance, while enabling population-level aggregation without storing multiple full sets of weights.
@@
6 self.z_instance = None # cached per sequence
@@
11 self.z = sample_normal(self.mu, self.sigma, shape=(B, 1, D), device=x.device)
@@
15 """Resamples the model instance (e.g., for a new prompt)."""
@@
20 for each normalization module to target: # locate layer by model.named_modules
21 replace it with BayesianNorm(module, mu, sigma)
@@
Figure 1: B-Trans implementation (abstracted). B-Trans operates as a plug-and-play replacement for normalization layers. For each response, it caches a latent offset per sequence and reuse it across autoregressive steps. This approximates drawing a single coherent model instance for each generation pass, enabling diverse yet temporally consistent outputs from one set of weights.
@@ 4 Implementation
Our approach is a surgical architectural modification rather than an external sampling procedure. We align with the pragmatist view that partial Bayesian structure in carefully chosen components can already induce useful functional properties (kristiadi2020being), we implement B-Trans by wrapping RMSNorm layers with a Bayesian post-normalization offset. The modified backbone remains differentiable and compatible with standard training pipelines. B-Trans operates as a plug-and-play replacement for normalization layers, ensuring that the backbone differentiable and compatible with standard training pipelines. Consequently, our method can be seamlessly integrated with existing training workflows.
@@ 4.1 Bayesian Bias Norm Wrapper
Our decision to target bias-like offsets is motivated by prior findings elhage2021mathematical; dong2021attention that biases encode “default” priors and are crucial for rank restoration, allowing us to shift reasoning dispositions without corrupting the semantic correlations stored in weights (see more discussion in Section 4.2). Concretely, for a hidden state $\mathbf{x}$, we implement:
$\mathbf{y}=\text{Norm}(\mathbf{x})\cdot\mathbf{w}+(\mathbf{b}+\mathbf{z}),\quad\mathbf{z}\sim\mathcal{N}(\mu,\sigma^{2})$ (4)
Here, $\mu$ and $\sigma$ are hyperparameters controlling the center and spread of the population diversity, and $\mathbf{b}$ can be 0 in RMSNorm. For normalization variants that include a learned scale (and optionally a bias), these parameters remain part of the deterministic backbone, while $\mathbf{z}$ provides a stochastic post-normalization offset. Sampling $\mathbf{z}$ therefore yields multiple model instances that share the same pre-trained weights but differ in their normalization regime.
4.2 Biases as Dispositional Priors.
B-Trans introduces stochasticity through additive offsets rather than through full weight perturbations. This choice is grounded in the functional role of additive shifts in Transformer mechanics. First, additive components act as input-independent defaults: while weight matrices encode correlations between features, additive shifts set baseline writes to the residual stream and influence activation thresholds before input-specific computation elhage2021mathematical. Second, additive vectors are a practical handle for semantic steering. Work on activation engineering shows that adding vectors in the residual stream can steer high-level behaviors without broadly degrading capabilities turner2023steering; zou2023representation. Third, bias-like shifts can be important for defining operating points. Recent evidence on massive activations suggests that when explicit biases are absent, models may simulate similar effects through outlier activations sun2024massive. By sampling $\mathbf{z}$, B-Trans creates a population of dispositions that explore different reasoning trajectories while sharing the same underlying knowledge in the backbone weights.

### v2

# Many Minds from One Model: Bayesian-Inspired Transformers for Population Diversity

@@ Abstract
Despite their scale and success, modern transformers are usually trained as single-minded systems: optimization produces a deterministic set of parameters, representing a single functional hypothesis about the data. Motivated by the analogy to human populations, in which population-level intelligence emerges from diverse individual behaviors, we propose Population Bayesian Transformers (B-Trans), which enable sampling diverse yet coherent transformer large language model instances (hereafter referred to as a ‘mind’) from a single pre-trained LLM.
B-Trans introduces a Bayesian-inspired posterior proxy by injecting stochasticity directly into normalization layers, avoiding the prohibitive cost of training full Bayesian neural networks. Sampling from this proxy yields a population of minds with diverse behaviors while maintaining general competence. During the generation of each response, we sample a single realization from the random distribution and hold it fixed, ensuring temporal consistency and reasoning coherence. Experiments on zero-shot generation and Reinforcement Learning with Verifiable Rewards (RLVR) demonstrate that B-Trans effectively leverages the stochastic model diversity, yielding superior response diversity while achieving better task performance compared to deterministic baselines.
@@
Bayesian-Inspired Transformers for Population Diversity
@@ 1 Introduction
The predominant paradigm in Large Language Model (LLM) Yang et al. (2025); Liu et al. (2025) deployment treats model weights as static, deterministic point estimates. While this approach effectively leverages the knowledge acquired during pre-training Wang et al. (2024), it inherently limits the model’s ability to generate diverse content, i.e., mode collapse Kirk et al. (2023). Recent studies further demonstrate that well-tuned models suffer from severe structural homogenization (Jiang et al., 2025). Under the convergence pressures of post-training alignment, such as Supervised Fine-Tuning (SFT) and Reinforcement Learning from Human Feedback (RLHF), models tend to collapse (converge) onto a narrow set of reasoning patterns. As a result, the model struggles to explore alternative reasoning paths, locking itself into local optima Yue et al. (2025). This rigidity, characterized by strong exploitation but limited exploration, can be viewed as a fundamental bottleneck in complex problem solving and learning for Artificial Intelligence Sutton et al. (1998).
@@
From the perspective of human intelligence, a population of decentralized minds can explore more insightful ideas than any single individual, a phenomenon known as the wisdom of crowds Surowiecki (2005). However, replicating this with a group of models is non-trivial, limited not only by high computational costs but also by the structural homogenization across model instances Jiang et al. (2025). Current attempts to reintroduce diversity at inference time operate mostly through injecting randomness into the action distribution, for example, through temperature-based sampling and related entropy-increasing strategies Wang et al. (2022). While increasing temperature can prevent verbatim repetition, it often acts as a surface-level perturbation. As noted by Yue et al. (2025), in reasoning tasks, sampling can lead to incoherent degeneration, yielding statistical variation without meaningful semantic progress, rather than systematically discovering novel solution paths. The core issue is that manipulating the output tokens does not alter the underlying deterministic reasoning engine.
@@
Can we achieve meaningful population diversity without abandoning the efficiency of a single model? In this work, we demonstrate the feasibility by generating many model instances by modifying a single pre-trained model. Our key insight is that normalization layers Ba et al. (2016); Zhang and Sennrich (2019) modulate representation geometry, and that small shifts, e.g., bias-like offsets, can induce diverse behaviors while preserving semantic meaningfulness. We therefore replace point-estimated normalization biases with Gaussian-distributed stochastic variables, transforming a single transformer to act as a population of model instances. Each LLM drawn from this distribution corresponds to a distinct mind [^1] In this paper, we use the term mind interchangeably with sampled model instance. We adopt this terminology to emphasize that each instance acts as a distinct individual from a population. that shares the same weights and structure, yet processes information through a different normalization parameter (i.e. due to different sampled offsets). This targeted population modeling yields diversity where it is influential, without incurring the cost of maintaining multiple full models.
@@
We present B-Trans as a general method for sampling populations of transformer-based large language models, where each sample represents an individual mind (Section 3). In this paper, we develop B-Trans through normalization-layer modification (Section 4), which is deliberately minimal and can be extended to other kinds of modifications in future work. To maintain logical coherence, we maintain temporal consistency by using the same sampled LLM when generating a response sequence. Unlike per-token noise injection (e.g., standard Dropout), which can lead to disjointed logic, our approach ensures that each response is produced by a single, logically consistent model instance sampled from the population, without changing the mind in the middle of a response.
@@
B-Trans is validated through two distinct experimental lenses, each demonstrating different benefits of population diversity. In zero-shot generation, we demonstrate that B-Trans effectively breaks structural homogenization, producing semantically meaningful, heterogeneous outputs. In Reinforcement Learning with Verifiable Rewards (RLVR) Guo et al. (2025), we show that the population-based B-Trans enables deep exploration, traversing sparse reward landscapes more effectively than deterministic model baselines. Collectively, these results confirm that B-Trans alters the exploratory dynamics of the model, enabling deep and population-driven reasoning at negligible computational cost.
@@ 2 Related Works
Bayesian methods provide a principled framework for modeling weight uncertainty (Blundell et al., 2015). Prior work has adapted Bayesian principles to neural networks for specific pragmatic objectives. For instance, Kristiadi et al. (2020) demonstrated that being “a bit Bayesian” (e.g., applying Laplace approximation only to the last layer) can improve calibration and reduce overconfidence (“knowing what you don’t know”). Similarly, Jing et al. (2025) introduced the Kalman Bayesian Transformer to support stable sequential learning and mitigate catastrophic forgetting by treating pre-trained weights as priors. While these works primarily use Bayesian tools for calibration, robustness, uncertainty quantification, or model averaging for prediction quality, B-Trans targets a generation-centric goal: we use lightweight parameter-space stochasticity to promote exploration and semantic diversity during generation. Compared to Deep Ensembles (Lakshminarayanan et al., 2017), which scale linearly in cost, or full BNNs, which are computationally intractable, B-Trans induces this population diversity solely through additive modification of normalization computing, offering a negligible alternative specifically optimized for modern LLMs.
@@
Standard decoding strategies, such as Temperature Ficler and Goldberg (2017), Top-$k$ Fan et al. (2018), and Nucleus Sampling Holtzman et al. (2019), primarily manipulate the token distribution to include diversity in model output. For example, methods that aggregate responses sampled via these strategies Wang et al. (2022) can provide slightly different outputs but still remain constrained by intra-mode collapse Kirk et al. (2023), i.e., a single model generates outputs with low semantic diversity due to the point-estimation nature of the underlying model.
Another line of research improves the answer diversity in some tasks by ensembling responses from multiple LLMs (population) at a considerable computational cost Jiang et al. (2023). However, recent work points out that modern instruction-tuned models exhibit severe inter-model homogeneity Jiang et al. (2025), where different LLMs independently converge on similar ideas with minor variations in phrasing.
B-Trans instead diversifies behavior by sampling model instances via parameter-space modification, turning a single set of weights into a population that can produce diverse generations.
@@
Balancing the tradeoff between exploration and exploitation during learning is a long-standing challenge for learning Sutton et al. (1998). Although huge efforts have been made by action-space entropy regularization during training Haarnoja et al. (2018); Schulman et al. (2017), in practice, Yue et al. (2025) observe that existing RL methods often force the model into incoherent degeneration (statistical noise) rather than systematically discovering alternative solution paths. Meanwhile, as aforementioned, empirical analysis also reveals the structural homogenization of well-aligned models, where the probability mass over-concentrates on a narrow set of reasoning patterns Jiang et al. (2025).
Alternatively, instead of addressing this intractable optimization problem, we attempt to improve output diversity during the decoding stage. We revisit the concept of Parameter Space Noise Plappert et al. (2018), originally proposed for continuous control, and adapt it to modern LLMs. B-Trans keeps the pretrained model weight and applies modifications to a lightweight posterior proxy during model inference. The goal is to achieve better exploration via diverse reasoning trajectories.
3 Bayesian-Inspired Population Model
This section formalizes a population view of a single pre-trained model via Bayesian hypothesis sampling: we draw a sequence-level parameter sample to obtain one coherent model instance, and aggregate multiple individual models for population-level thinking. [^2] Throughout, “Bayesian” is used as a principled motivation for structured uncertainty and sampling, rather than a claim of exact posterior inference for billion-parameter LLMs. Section 4 then presents a minimal instantiation that perturbs normalization offsets as a computationally negligible proxy for posterior sampling.
@@ 3.1 Local Geometry and the Laplace Approximation
Let $\mathcal{D}$ be the training dataset, and $\theta\in\mathbb{R}^{d}$ be the parameters of a Transformer. We view the pre-trained parameters $\theta_{\text{MAP}}$ as the Maximum A Posteriori (MAP) estimate of the posterior $p(\theta|\mathcal{D})$. While $\theta_{\text{MAP}}$ captures a high-probability configuration, it collapses the rich information contained in the geometry of the loss landscape into a single point. To recover uncertainty, we employ the Laplace Approximation, modeling the posterior proxy around the mode as a Gaussian:
@@
Computing the exact covariance $\Sigma$ (the inverse Hessian) is not feasible for billion-parameter models. Furthermore, the true posterior landscape of LLMs is highly non-isotropic (Li et al., 2018), and estimating its anisotropy at inference time without access to training data or gradients is not well-posed.
In principle, constructing the exact posterior geometry presents two pathways: (1) performing full Bayesian training to learn the parameter distribution from vast data associated with different LLMs (text-author pairs), or (2) computing the inverse Hessian matrix. However, both approaches are computationally prohibitive and data-intensive for modern LLMs. We therefore opt for a tractable approximation by empirically defining an isotropic covariance structure $\Sigma=\sigma^{2}I$. From an information-theoretic perspective, this choice represents the Maximum Entropy distribution given a constraint on the expected Euclidean distance from the $\theta_{\text{MAP}}$ estimate. Rather than risking a mis-specified diagonal approximation that might bias the model into degenerate subspaces, the isotropic assumption effectively defines a spherical trust region for exploration. Importantly, our goal is not exact Bayesian inference, but efficient sampling from a locally valid posterior proxy, and a practical balancing of theoretical motivation with engineering feasibility. While this variance is treated as a structural hyperparameter rather than a learned parameter, we show that this simplification is empirically sufficient to induce meaningful population diversity without the intractable costs of the full Bayesian alternative
@@
A key challenge for Bayesian sampling in autoregressive generation is handling the time dimension. Standard stochastic regularizers, such as Dropout (Srivastava et al., 2014), are typically applied independently at each forward pass. For a sequence $y_{1:T}$, this implies integrating over the parameters at every token step:
@@
We argue that Eq. (2) is misaligned with coherent multi-step reasoning: it re-samples $\theta_{t}$ at each token, which can disrupt cross-token logical consistency by changing the underlying computation mid-generation. Instead, we adopt hypothesis sampling. In the Bayesian view, a single sample $\theta\sim p(\theta|\mathcal{D})$ corresponds to one coherent hypothesis. Once sampled, this hypothesis should remain fixed for the entire sequence. This also fits the real-world scenario, as each sequence is normally written by one person (which corresponds to one LLM in our sample). The predictive distribution is:
@@
We approximate Eq. (3) using Monte Carlo integration with $K$ samples. For each sample $k$, we draw a noise vector $\epsilon^{(k)}$ once and freeze it, obtaining a specific model member $\theta^{(k)}=\theta_{\text{MAP}}+\epsilon^{(k)}$. This ensures that the entire sequence $y_{1:T}$ is generated by a single consistent LLM instance, while enabling population-level aggregation without storing multiple full sets of weights.
@@
6 self.z = None # cached per sequence
@@
11 self.z = sample_normal(self.mu, self.sigma, shape=(B, 1, D))
@@
15 """Resamples the model instance (e.g., for a new mind)."""
@@
20 for each normalization layer: # locate layer by model.named_modules
21 replace it with BayesianBiasWrapper(module, mu, sigma)
@@
Figure 1: B-Trans implementation (abstracted). B-Trans operates as a plug-and-play replacement for normalization layers. For each response, it caches a latent offset per sequence and reuses it across autoregressive steps. This approximates drawing a single coherent model instance for each generation pass, enabling diverse yet temporally consistent outputs from one set of weights.
@@ 4 Implementation
Our approach is a surgical architectural modification during the model inference phase. Aligning with the pragmatist view that partial Bayesian structure in carefully chosen components can already induce useful responses (Kristiadi et al., 2020), we implement B-Trans by wrapping RMSNorm layers with a sampled offset. B-Trans operates as a plug-and-play replacement for normalization layers, ensuring that the modified LLM is still compatible with existing inference engines such as vLLM Kwon et al. (2023). Furthermore, B-Trans remains differentiable, which makes backpropagation possible for all kinds of learning in future works.
@@ 4.1 Bayesian Bias Norm Wrapper
Our decision to target bias-like offsets is motivated by prior findings Elhage et al. (2021); Dong et al. (2021) that biases encode “default” priors and are crucial for rank restoration, allowing us to shift reasoning dispositions without corrupting the semantic correlations stored in weights (see more discussion in Section 4.2). Concretely, for a hidden state $\mathbf{x}$, we implement:
$\mathbf{y}=\mathrm{Norm}(\mathbf{x};\mathbf{w},\mathbf{b})+\mathbf{z},\quad\mathbf{z}\sim\mathcal{N}(\boldsymbol{\mu},\,\sigma^{2})$ (4)
Here, $\mu$ and $\sigma$ are hyperparameters controlling the center and spread of the population diversity, and $\mathbf{b}$ can be 0 in RMSNorm (which makes $z$ a post-norm additive offset). For normalization variants that include a learned scale (and optionally a bias), these parameters remain part of the deterministic backbone, while $\mathbf{z}$ provides a stochastic post-normalization offset for each sampled model instance.
4.2 Biases as Dispositional Priors
B-Trans introduces stochasticity through additive offsets rather than through full weight modification. This choice is grounded in the functional role of additive shifts in Transformer mechanics. First, additive components act as input-independent defaults: while weight matrices encode correlations between features, additive shifts set baseline writes to the residual stream and influence activation thresholds before input-specific computation Elhage et al. (2021). Perturbing these terms, therefore, alters a model’s inductive dispositions without corrupting the semantic relationships stored in its weights. Second, additive vectors are a practical handle for semantic steering. Works on activation engineering show that adding vectors in the residual stream can steer high-level behaviors without broadly degrading capabilities Turner et al. (2023); Zou et al. (2023). Third, bias-like shifts can be critical for model behavior. Sun et al. (2024) suggests that when explicit biases are absent, models may simulate similar effects through outlier activations. For these reasons, B-Trans targets additive offsets as a minimal and controlled locus for inducing population diversity. While extending stochasticity to additional parameters—such as normalization weights $w$ (Equation 4)—may further enrich diversity, we do not explore such extensions in this work. Perturbing full weight matrices of LLMs, in contrast, risks disrupting learned representations and introducing excessive noise, potentially degrading model behavior. We therefore restrict our study to additive offsets and leave broader normalization- or weight-level stochasticity to future work.